OpenStudy (anonymous):

Solve the system using elimination. 4x + 2y = –10 4x – 2y = –14

OpenStudy (anonymous):

I would add them resulting in: 4x + 2y + 4x -2y = -10 - 14 8x = -24 x = -3 I think you can solve the rest :)

OpenStudy (anonymous):

4x+2y=-10 4x+2y=-14 8x=-24 divide by 8 x=-3 4x+2y=-10 -(4x+2y=-14)= -4x-2y=14 y=4